OpenStudy (anonymous):

Melania has two jobs. During the day, she works as an office clerk, and in the evening she works as a cashier. Her office job pays her $13 per hour, and her cashier job pays her $9.25 per hour. In one week, Melania worked 38 hours, and she earned a total of $434. If c represents the number of hours that Melania works as an office clerk, write an equation that can be used to find how many hours Melania worked in both jobs. (Points : 1) 13c + 9.25(c – 38) = 434 13c + 9.25(38 – c) = 434 13c + 9.25(434 – c) = 38 9.25c + 13(38 – c) = 434

OpenStudy (anonymous):

@Algorithmic please help me please i have this than one more and then I'm done

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Glen decides he only wants caramels and candy corn. Caramels sell for $0.95 per pound, and candy corn sells for $0.79 per pound. Glen’s bag weighs 1.7 pounds and it costs $1.48. If p = pounds of caramels, which equation in simplest form represents the situation described. (Points : 1) –0.16p + 1.62 = 1.48 –0.16p + 1.62= 1.7 0.16p + 1.34 = 1.48 0.16p + 1.34 = 1.7

OpenStudy (anonymous):

sorry I'm this one now

OpenStudy (anonymous):

please just help me fast

OpenStudy (kkutie7):

\[caramels=\frac{0.95dollars}{lbs}\] \[candycorn=\frac{0.79}{lbs}\] \[total=\frac{1.48dollars}{1.7lbs}\]

OpenStudy (anonymous):

its 2 o'clock and im so so tired and frustrated and sad

OpenStudy (anonymous):

sorry for rushing your so great thank you

OpenStudy (kkutie7):

\[\frac{1.48dollars}{1.7lbs}=\frac{.79dollars}{lbs}+\frac{.95dollars}{?lbs}\]

OpenStudy (anonymous):

C? i was thinkikng

OpenStudy (anonymous):

@Kkutie7

OpenStudy (kkutie7):

yeah its either A or C

OpenStudy (anonymous):

A store allows customers to fill their own bags of candy. Troy decides he only wants chocolate-covered pretzels and gumdrops. Chocolate-covered pretzels sell for $0.89 per pound, and gumdrops sell for $0.65 per pound. Troy’s bag weighs 1.8 pounds and it cost $1.29. How many pounds of chocolate-covered pretzels and gumdrops did Troy buy? (Points : 1) 0.5 pounds of pretzels; 1.3 pounds of gumdrops 0.9 pounds of pretzels; 0.9 pounds of gumdrops 1.3 pounds of pretzels; 0.5 pounds of gumdrops 0.8 pounds of pretzels; 1 pound of gumdrops

OpenStudy (anonymous):

thanks by the way